OpenStudy (anonymous):

how do you solve b/16=b-6/4 b over 16=b-6 over 4 is how it is written in book

hero (hero):

Hint: Cross multiply

OpenStudy (anonymous):

b=16x(b-6\4) ; cross multiplication b=16b-(16x6\4); we know 4x4=16 implies 16\4=4 b=16b-4x6 16b-b=24 15b=24 therefore b=24\15 clear?